### v1

# Spherical solutions to the Klein-Gordon equation in the expanding universe

Abstract
We produce an explicit formula for the wave function of the spherically symmetric fields emitted to the FLRW universe with the scale factor generated by the de Sitter universe. As an application of these explicitly written solutions of the Klein-Gordon equation, we test the decay in time of the field generated by a pionic atom.
Keywords: Klein-Gordon equation, Spherical waves, Expanding universe, Pionic atom
School of Mathematical and Statistical Sciences,
University of Texas RGV, 1201 W. University Drive,
Edinburg, TX 78539, USA
1 Introduction. Main Results
In this study, we produce an explicit wave function of the particles (fields) emitted to the FLRW universe with the scale factor generated by the de Sitter universe. We consider models of scalar fields with spherical symmetry at the initial time. They can be generated, for instance, by a pionic atom, exotic atom, or gravitational field (see, e.g., [8, 14, 5, 26]). In the expanding universe, the particles obey the equation for the free particles. A purpose of this article is a derivation of the explicit form of exact solutions of the spherical waves in the curved spacetime. The models illustrate an instantaneous transition from the initial spacetime to the non-static FLRW universe with a de Sitter scale factor.
Thus, for the scalar field, we are interested in the Cauchy problem
$\cases{\displaystyle\Box_{g}\Phi+\frac{m_{n}^{2}c^{2}}{\hbar^{2}}\Phi=0\,,\cr\displaystyle\Phi(x,0)=F(r,0)Y_{\ell m}(\theta,\phi)\,,\quad\Phi_{t}(x,0)=F_{t}(r,0)Y_{\ell m}(\theta,\phi)\,,}$
where $m_{n}$ is a mass, $F(r,t)$ is a time-dependent radial part of the initial wave, $F_{t}(r,t):={\partial_{t}F}(r,t)$, while $Y_{\ell m}(\theta,\phi)$ are spherical harmonics. The line element of the non-stationary spacetime with the metric tensor $g_{ik}$ and the scale factor $a=a(t)$ in the spherical coordinates is
$ds^{2}=-c^{2}dt^{2}+a^{2}(t)dr^{2}+a^{2}(t)r^{2}(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}).$ (1.1)
For the de Sitter universe in FLRW form, $a(t)=e^{Ht}$, where $H$ is the Hubble parameter. The scalar curvature of this space is $12H^{2}$. Consequently, the covariant d’Alembert’s operator in the spacetime with the metric of (1.1) is
$\square_{g}\Phi=\frac{1}{\sqrt{|g|}}\frac{\partial}{\partial x^{i}}\left(\sqrt{|g|}g^{ik}\frac{\partial\Phi}{\partial x^{k}}\right).$
We are interested in the explicit form of the spherical field in the de Sitter FLRW spacetime, which is a solution of the problem
$\cases{\displaystyle\Phi_{tt}+3H\Phi_{t}-c^{2}e^{-2Ht}\Delta\Phi+\frac{m_{n}^{2}c^{4}}{\hbar^{2}}\Phi=0\,,\cr\displaystyle\Phi(x,0)=F_{0}(r)Y_{\ell m}(\theta,\phi)\,,\quad\Phi_{t}(x,0)=F_{1}(r)Y_{\ell m}(\theta,\phi)\,.}$ (1.2)
Here, $Y_{lm}(\theta,{\phi})$ is a spherical harmonic, $\ell=0,1,\ldots$, $m=0,\pm 1,\pm 2\ldots$, $|m|\leq\ell$. In particular, the initial state can be given by a wave function $\mbox{\rm Exp}\left(-i\frac{E_{n}t}{\hbar}\right)\frac{R_{n\ell}(r)}{r}Y_{\ell m}(\theta,\phi)$ of the pionic atom (see, e.g., [8]). We will use a system of units in which the speed of light $c=1$.
Denote $M:=\left(\frac{9}{4}H^{2}-\frac{m_{n}^{2}}{\hbar^{2}}\right)^{1/2}$. Let $F(a,b;c;z)$ be a hypergeometric function (see, e.g., [3]), and $J_{\nu}(z)$ be a Bessel function. We define kernels $K_{0}(z,t;M)$ and $K_{1}(z,t;M)$ as  [^1] There is a typo in $K_{0}$ on page 11 [26]; the correct version is in Lemma 3.1 [26]
$K_{0}(r,t;M)$ $:=$ $-4^{-\frac{M}{H}}e^{-Ht}e^{Mt}\left(\left(e^{-Ht}+1\right)^{2}-H^{2}r^{2}\right)^{\frac{M}{H}-\frac{5}{2}}$
$\times\Bigg{\{}e^{Ht}\left(\left(e^{-Ht}+1\right)^{2}-H^{2}r^{2}\right)\left(-H^{2}Mr^{2}+Me^{-2Ht}+He^{-Ht}+H-M\right)$
$\times F\left(\frac{1}{2}-\frac{M}{H},\frac{1}{2}-\frac{M}{H};1;\frac{\left(1-e^{-Ht}\right)^{2}-H^{2}r^{2}}{\left(1+e^{-Ht}\right)^{2}-H^{2}r^{2}}\right)$
$+\frac{1}{H}(H-2M)^{2}\left(-H^{2}r^{2}+e^{-2Ht}-1\right)$
$\times F\left(\frac{3}{2}-\frac{M}{H},\frac{3}{2}-\frac{M}{H};2;\frac{\left(1-e^{-Ht}\right)^{2}-H^{2}r^{2}}{\left(1+e^{-Ht}\right)^{2}-H^{2}r^{2}}\right)\Bigg{\}},$
$K_{1}(r,t,M)$ $:=$ $4^{-\frac{M}{H}}e^{Mt}\left(\left(e^{-Ht}+1\right)^{2}-H^{2}r^{2}\right)^{\frac{M}{H}-\frac{1}{2}}$
$\times F\left(\frac{1}{2}-\frac{M}{H},\frac{1}{2}-\frac{M}{H};1;\frac{\left(1-e^{-Ht}\right)^{2}-H^{2}r^{2}}{\left(1+e^{-Ht}\right)^{2}-H^{2}r^{2}}\right)\,.$
We often omit the parameter $M$ in $K_{0}(z,t;M)$ and $K_{1}(z,t;M)$ when it does not lead to confusion. Denote $\phi(t):=(1-e^{-Ht})/H$. The main results of this study, which are obtained by the integral transform approach (ITA) [21, 24], are the following two theorems.
Theorem 1.1
Assume that $F_{0}(r),F_{1}(r)\in C^{1}({\mathbb{R}}\setminus\{0\})$, $\ell$ is non-negative integer number, and the condition
$F_{i}(r)=O(r^{\mu-\frac{1}{2}})\,\,\mbox{\rm as }\,\,r\searrow 0,\quad i=0,1\,$,
is fulfilled with $\mu>\ell-\frac{3}{2}$. Further, assume that $r^{\ell}F_{0}(r)$ and $r^{\ell}F_{1}(r)$ are even functions. If $Y_{lm}(\theta,{\phi})$ is a spherical harmonic, $\ell=0,1,\ldots$, $m=0,\pm 1\pm 2,\ldots$, and $|m|\leq\ell$, then the spherical field in the de Sitter FLRW spacetime is given as follows
$\Phi(r,\theta,\phi,t)$
$=$ $e^{-Ht}Y_{\ell m}(\theta,\phi)\Bigg{\{}\frac{1}{2r}[(r-\phi(t))F_{0}(r-\phi(t))+(r+\phi(t))F_{0}(r+\phi(t))]$
$-\frac{1}{4}\ell(\ell+1)\frac{\phi(t)}{r^{2}}\int_{r-\phi(t)}^{r+\phi(t)}F_{0}(s)F\left(1-\ell,\ell+2;2;\frac{\phi(t)^{2}-(r-s)^{2}}{4rs}\right)ds\Bigg{\}}$
$+\,e^{-H\frac{3}{2}t}Y_{\ell m}(\theta,\phi)\int_{0}^{\phi(t)}\Bigg{\{}\frac{1}{2r}[(r-\tau)F_{0}(r-\tau)+(r+\tau)F_{0}(r+\tau)]$
$-\frac{1}{4}\ell(\ell+1)\frac{\tau}{r^{2}}\int_{r-\tau}^{r+\tau}F_{0}(s)F\left(1-\ell,\ell+2;2;\frac{t^{2}-(r-s)^{2}}{4rs}\right)ds\Bigg{\}}\left(2K_{0}\left(\tau,t\right)+3K_{1}\left(\tau,t\right)\right)\,d\tau$
+ $+\,2e^{-\frac{3}{2}Ht}Y_{\ell m}(\theta,\phi)\int_{0}^{\phi(t)}\Bigg{\{}\frac{1}{2r}[(r-\tau)F_{1}(r-\tau)+(r+\tau)F_{1}(r+\tau)]$
$-\frac{1}{4}\ell(\ell+1)\frac{\tau}{r^{2}}\int_{r-\tau}^{r+\tau}F_{1}(s)F\left(1-\ell,\ell+2;2;\frac{\tau^{2}-(r-s)^{2}}{4rs}\right)ds\Bigg{\}}K_{1}\left(\tau,t\right)\,d\tau\,.$
For the huygensian scalar field (particle) (1.2), that is, $m_{n}^{2}=2H^{2}\hbar^{2}$, the spherical field in the de Sitter FLRW spacetime is
$\Phi(r,\theta,\phi,t)$
$=$ $e^{-Ht}Y_{\ell m}(\theta,\phi)\Bigg{\{}\frac{1}{2r}[(r-\phi(t))F_{0}(r-\phi(t))+(r+\phi(t))F_{0}(r+\phi(t))]$
$-\frac{1}{4}\ell(\ell+1)\frac{\phi(t)}{r^{2}}\int_{r-\phi(t)}^{r+\phi(t)}F_{0}(s)F\left(1-\ell,\ell+2;2;\frac{\phi(t)^{2}-(r-s)^{2}}{4rs}\right)ds\Bigg{\}}$
$+\,e^{-Ht}Y_{\ell m}(\theta,\phi)\int_{0}^{\phi(t)}\Bigg{\{}\frac{1}{2r}\Big{(}(r-\tau)[F_{0}(r-\tau)+F_{1}(r-\tau)]+(r+\tau)[F_{0}(r+\tau)+F_{1}(r+\tau)]\Big{)}$
$-\frac{1}{4}\ell(\ell+1)\frac{\tau}{r^{2}}\int_{r-\tau}^{r+\tau}(F_{0}(s)+F_{1}(s))F\left(1-\ell,\ell+2;2;\frac{\tau^{2}-(r-s)^{2}}{4rs}\right)ds\Bigg{\}}d\tau\,.$
It should be noted here that for negative integers $a$, the hypergeometric function $F(a,b;c;z)$ is a polynomial of degree a-a (see, e.g., [3]).
In the next theorem, by imposing conditions on the behavior of the functions $F_{0}=F_{0}(r)$ and $F_{1}=F_{1}(r)$ for large $r$ one can appeal to the Hankel transform and write the representation of the field in the form that is more amendable for the $L^{2}({\mathbb{R}}^{3})$-estimates. In particular, this applies to the wave function of the pionic atom that instantly arrives in the de Sitter FLRW spacetime, because that wave function has an exponential decay at infinity of spatial coordinates (see, e.g., [8]).
Theorem 1.2
Assume that $F_{0}(r),F_{1}(r)\in C^{1}(0,\infty)$, and
$\int_{0}^{\infty}r^{3/2}|F_{i}(r)|\,dr<\infty,\quad F_{i}(r)=O(r^{-k}),\quad r\to\infty,\quad k>2,\quad i=0,1.$
If $Y_{lm}(\theta,{\phi})$ is a spherical harmonic, $\ell=0,1,\ldots$, $|m|\leq\ell$, then the spherical field in the de Sitter FLRW spacetime with the initial state
$\Phi(x,0)=F_{0}(r)Y_{\ell m}(\theta,\phi)\,,\quad\Phi_{t}(x,0)=F_{1}(r)Y_{\ell m}(\theta,\phi)$
is given by
$\Phi(r,\theta,\phi,t)$
$=$ $e^{-Ht}Y_{\ell m}(\theta,\phi)\frac{1}{\sqrt{r}}\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{0}(\rho)J_{\ell+\frac{1}{2}}(\rho\lambda)\rho^{3/2}\,d\rho\right)\cos\left(\lambda\phi(t)\right)J_{\ell+\frac{1}{2}}(r\lambda)\lambda\,d\lambda$
$+\,e^{-\frac{3}{2}Ht}Y_{\ell m}(\theta,\phi)\frac{1}{\sqrt{r}}\int_{0}^{1}\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{0}(\rho)J_{\ell+\frac{1}{2}}(\rho\lambda)\rho^{3/2}\,d\rho\right)\cos\left(\lambda\phi(t)s\right)$
$\times J_{\ell+\frac{1}{2}}(r\lambda)\lambda\,d\lambda\left(2K_{0}\left(\phi(t)s,t\right)+3K_{1}\left(\phi(t)s,t\right)\right)\phi(t)\,ds$
+ $+\,2e^{-\frac{3}{2}Ht}Y_{\ell m}(\theta,\phi)\frac{1}{\sqrt{r}}\int_{0}^{1}\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{1}(\rho)J_{\ell+\frac{1}{2}}(\rho\lambda)\rho^{3/2}\,d\rho\right)\cos\left(\lambda\phi(t)s\right)$
$\times J_{\ell+\frac{1}{2}}(r\lambda)\lambda\,d\lambda\,K_{1}\left(\phi(t)s,t\right)\phi(t)\,ds\quad\mbox{ for all}\quad x\in{\mathbb{R}}^{3},\,\,t>0.$
For the huygensian scalar field (particle) (1.2), that is, $m_{n}^{2}=2H^{2}\hbar^{2}$, the spherical field in the de Sitter FLRW spacetime is
$\Phi(r,\theta,\phi,t)$
$=$ $e^{-Ht}Y_{\ell m}(\theta,\phi)\frac{1}{\sqrt{r}}\Bigg{\{}\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{0}(\rho)J_{\ell+\frac{1}{2}}(\rho\lambda)\rho^{3/2}\,d\rho\right)\cos\left(\lambda\phi(t)\right)J_{\ell+\frac{1}{2}}(r\lambda)\lambda\,d\lambda$
$+\int_{0}^{\infty}\left(\int_{0}^{\infty}(F_{0}(\rho)+F_{1}(\rho))J_{\ell+\frac{1}{2}}(\lambda\rho)\,\rho^{3/2}d\rho\right)\sin\left(\lambda\phi(t)\right)J_{\ell+\frac{1}{2}}(r\lambda)\,d\lambda\Bigg{\}}\,.$
Below, we provide some examples of such an initial state $(\Phi(x,0),\Phi_{t}(x,0))$ of the field. In particular, it can be a wave function of a pionic atom in the Minkowski spacetime (see, e.g., [8, 10]), or a scalar field in the matter- or radiation-dominated or de Sitter universe with a different Hubble constant $H_{0}\not=H$ and different mass. In exotic atoms, one of the electrons has been replaced by a negatively charged particle such as a muon, K-meson, pion, $\Sigma$-hyperon, etc. For further details on the physical interest in such systems, see, for example, [8, 9, 10, 13]. A pionic atom is a hydrogen-like system in which an electron replaced by a spinless relativistic particle pion. A pion is an example of a particle with electromagnetic interactions that obeys the Klein-Gordon equation. In fact, by pionic atom, one can test the validity of the Klein-Gordon equation (see, e.g., [10, Sec.18.4]). In (1.2), the pion has instantly left the atom, the Coulomb field is no longer in action, and spacetime has instantly become inflationary. Thus, the free pion propagates in the expanding universe and carries information about its previous state. Another possible application can be to the problem of the existence of quasinormal modes on de Sitter space (see [12]).
Although spherically symmetric solutions for the Klein-Gordon and Dirac equations in the Minkowski spacetime have been known for a long time (see, e.g., [4, §24], [7], [15], and references therein), explicit exact formulas in the curved spacetimes are known only for a few special cases (see, e.g., [2, 6, 18, 25]). We believe that the obtained explicit formulas of Theorem 1.2 can be useful, in particular, in the study of caustics for spherical waves in curved spacetime (see, e.g., [7]). In the forthcoming paper, the representation formulas of Theorems 1.1, 1.2 will be used for the derivation of the exact spherical solutions to the spin-$\frac{1}{2}$ fields in the expanding universe.
The rest of this paper is organized as follows. In Section 2, we describe the explicit formulas for spherical solutions to the wave equation in the Minkowski space. We list three different approaches in subsections 2.1, 2.2, and 2.3, more exactly, approaches by the general solution, by the Riemann function, and by the Hankel transform, respectively. In Section 3 we apply the ITA to write the explicit form of the Klein-Gordon fields in the Minkowski and de Sitter spacetimes. This completes the proof of Theorems 1.1,1.2.
2 Spherical solutions of the wave equation in
Minkowski spacetime
To make the presentation of the proof of Theorem 1.2 self-contained, in this section, we write explicit solutions to the wave equation with the initial data, which are the products of the spherical harmonic and the radial function. More information on this type of result can be found, for instance, in the review [17], or in [1] and [20]. In our study, the radial function can be a radial part of the wave function of the free spinless particle, or pionic atom or can be generated by another model.
Consider the problem for the wave equation
$V_{tt}-\Delta V=0,\quad V(x,0)=V_{0}(x),\quad V_{t}(x,0)=V_{1}(x)\,.$ (2.1)
Here $\Delta$ is the Laplace operator in spherical coordinates
$\Delta=\partial_{r}^{2}+\frac{2}{r}\partial_{r}+\frac{1}{r^{2}}\partial_{\theta}^{2}+\frac{\cot(\theta)}{r^{2}}\partial_{\theta}+\frac{\csc^{2}(\theta)}{r^{2}}\partial_{\phi}^{2}=\partial_{r}^{2}+\frac{2}{r}\partial_{r}+\frac{1}{r^{2}}\Delta_{\theta,\phi}\,.$
The initial functions are assumed to be of the form
$V_{0}(x)=F_{0}(r)Y_{\ell m}(\theta,\phi),\quad V_{1}(x)=F_{1}(r)Y_{\ell m}(\theta,\phi),$ (2.2)
where $Y_{\ell m}(\theta,\phi)$, $|m|\leq\ell$, $\ell=0,1,\ldots$, are spherical harmonics,
$\Delta_{\theta,\phi}Y+\ell(\ell+1)Y=0,\quad\int\left|Y_{\ell m}(\theta,\phi)\right|^{2}d\Omega=1.$
The functions $F_{0}$ and $F_{1}$ are smooth in $\mathbb{R}$$_{+}$ and satisfy certain conditions as $r\to 0$ and $r\to\infty$. Consequently, we look for a solution of the form
$V(x,t)=F(r,t)Y_{\ell m}(\theta,\phi)\,.$
Then
$\cases{\displaystyle Y_{\ell m}(\theta,\phi)\frac{\partial^{2}}{\partial t^{2}}F(r,t)-Y_{\ell m}(\theta,\phi)\frac{1}{r}\frac{\partial^{2}}{\partial r^{2}}(rF(r,t))-F(r,t)\frac{1}{r^{2}}\Delta_{\theta,\phi}Y_{\ell m}(\theta,\phi)=0,\cr F(r,0)=F_{0}(r),\quad F_{t}(r,0)=F_{1}(r),\quad r\geq 0\,.}$
Hence,
$F_{tt}(r,t)-F_{rr}(r,t)-\frac{2}{r}F_{r}(r,t)+\frac{\ell(\ell+1)}{r^{2}}F(r,t)=0\,.$ (2.3)
For the pionic atom the wave function is
$\psi(x,t)=e^{itE}F(r,t)Y_{\ell m}(\theta,\phi)\,$,
where for the function $F=F(r,t)$ the radial equation is
$F_{tt}(r,t)-F_{rr}(r,t)-\frac{2}{r}F_{r}(r,t))+\left(U(r)+\frac{\ell(\ell+1)}{r^{2}}\right)F(r,t)=0,\quad r>0,\quad t\in{\mathbb{R}}\,$,
while $U$ is a potential. For the function $R(r,t)=rF(r,t)$, the equation is
$R_{tt}(r,t)-R_{rr}(r,t)+\left(U(r)+\frac{\ell(\ell+1)}{r^{2}}\right)R(r,t)=0,\quad r>0\,.$
The normalization conditions are
$\int_{0}^{\infty}R^{2}(r,t)\,dr<\infty,\qquad\int_{0}^{\infty}F^{2}(r,t)r^{2}\,dr<\infty.$ (2.4)
2.1 The general solution approach
Consider for the wave equation (2.1) the problem with the initial conditions
$V(x,0)=F_{0}(r)Y_{\ell m}(\theta,\phi),\quad V_{t}(x,0)=F_{1}(r)Y_{\ell m}(\theta,\phi)\,.$
We look for the solution $F=F(r,t)$ of equation (2.3) satisfying (2.4). For example, if the initial state $F_{0}(r)$ is the the wave function of the pionic atom [8, pp.188-189], [10, Sec.18.4], and [14, §26, Sec 2] in the Minkowski space, then
$F_{0}(r)=C_{n,\ell,\mu}r^{\mu-\frac{1}{2}}e^{-r/2}L_{n-\ell-1}^{2\mu}(r),\quad L_{n-\ell-1}^{2\mu}(r)\quad\mbox{\rm is Laguerre polynomial},\,\,r>0,$ (2.5)
where with the fine structure constant $\alpha=e^{2}/(\hbar c)\approx 1/137$ and the charge $e$ of the electron in Gaussian units,
$\mu=\sqrt{\left(\ell+\frac{1}{2}\right)^{2}-Z^{2}\alpha^{2}}\geq 0\,.$
If $\ell=0$, while $Z=1$ and $Z=2$, then $\mu-\frac{1}{2}=-0.0000532822$ and $\mu-\frac{1}{2}=-0.000213163$, respectively.
For the case of $F_{1}(r)=0$, assume that $F_{0}(r)$ is smooth on ${\mathbb{R}}\setminus\{0\}$. Define the operator
$L_{\ell}(r,\partial_{r},\partial_{t})$ $:=$ $\partial_{t}^{2}-\partial_{r}^{2}-\frac{2}{r}\partial_{r}+\frac{\ell(\ell+1)}{r^{2}}\,.$
It is easily seen that
$\left(\partial_{r}-\frac{\ell}{r}\right)L_{\ell}$ $=$ $L_{\ell+1}\left(\partial_{r}-\frac{\ell}{r}\right)\,.$
Hence, if $w_{(\ell)}=w_{(\ell)}(r,t)$ is a solution to $L_{\ell}(r,\partial_{r},\partial_{t})w_{(\ell)}(r,t)=0$, then the function
$w_{(\ell+1)}(r,t)=\left(\partial_{r}-\frac{\ell}{r}\right)w_{(\ell)}(r,t)$
solves the equation
$L_{\ell+1}(r,\partial_{r},\partial_{t})w_{(\ell+1)}(r,t)=0.$
Therefore, the general solution to equation (2.3) is given by the formula
$F^{(\ell)}(r,t)=r^{\ell}\left(\frac{1}{r}\frac{\partial}{\partial r}\right)^{\ell}\left(\frac{\Phi(r+t)+\Psi(r-t)}{r}\right),\quad r\in{\mathbb{R}},\quad t\geq 0\,,\quad\ell=0,1,2,\ldots\,$,
where $\Phi$ and $\Psi$ are the arbitrary functions. For every given $\ell$, one can derive from the general solution an explicit formula for the solution to the Cauchy problem (2.1)&(2.2), provided that (2.4) holds. The calculations lead to the following result. The solution $F^{(\ell)}(r,t)$, $\ell=0,1,2,\ldots$, of problem (2.3) with the initial data
$F^{(\ell)}(r,0)=F_{0}^{(\ell)}(r),\qquad F_{t}^{(\ell)}(r,0)=0,\qquad r\in{\mathbb{R}}_{+}\,$,
is given recursively by
$F^{(\ell+1)}(r,t)=\left(\partial_{r}-\frac{\ell}{r}\right)\widetilde{F}^{(\ell)}(r,t),\qquad\ell=0,1,2,\ldots\,$, (2.6)
where $\widetilde{F}^{(\ell)}(r,t)$ is given by
$\widetilde{F}^{(\ell)}(r,t)$ $=$ $\frac{1}{2r}\left\{(r-t)\widetilde{F}_{0}^{(\ell)}(r-t)+(r+t)\widetilde{F}_{0}^{(\ell)}(r+t)\right\}$
$+\frac{1}{2r^{\ell+1}}t\cases{\displaystyle\sum_{k\in{\mathbb{Z}}\,\,is\,\,odd,\,\,|k|\leq\ell-1\atop{\alpha+\beta+k=\ell-1\atop{\alpha\geq 0,\beta\geq 0\,\,are\,\,even}}\,\,}a_{k,\alpha,\beta}^{(\ell)}r^{\alpha}t^{\beta}\int_{r-t}^{r+t}\tau^{k}\widetilde{F}_{0}^{(\ell)}(\tau)\,d\tau,\quad\ell\,\,\mbox{is even},\cr{}\cr\displaystyle\sum_{\alpha+\beta+k=\ell-1\atop{\atop{k\in{\mathbb{Z}}\,\,is\,\,even,\,\,|k|\leq\ell-1\atop{\alpha\geq 0,\beta\geq 0\,\,are\,even}}}\,\,}b_{k,\alpha,\beta}^{(\ell)}r^{\alpha}t^{\beta}\int_{r-t}^{r+t}\tau^{k}\widetilde{F}_{0}^{(\ell)}(\tau)\,d\tau,\quad\ell\,\,\mbox{is odd}}$
and
$\widetilde{F}_{0}^{(\ell)}(r)=r^{\ell}\int_{0}^{r}s^{-\ell}F_{0}^{(\ell+1)}(s)\,ds\,.$ (2.8)
For the first three values of $\ell$, if $F_{0}$ is given by (2.5), then the function $F(r,t)$ can be a distributional solution from the Sobolev space. For the first five values of $\ell$, the results are as follows. For $\ell=0$ and the even function $F_{0}(r)$, the solution is
$F(r,t)$ $=$ $\frac{1}{2r}\left\{(r-t)F_{0}(r-t)+(r+t)F_{0}(r+t)\right\},\quad t\geq 0\,.$
For $\ell=1$ and odd function $F_{0}(r)$, the solution is
$F(r,t)$ $=$ $\frac{1}{2r}\left\{(r-t)F_{0}(r-t)+(r+t)F_{0}(r+t)\right\}-\frac{1}{2r^{2}}t\int_{r-t}^{r+t}F_{0}(\tau)\,d\tau\,.$
For $\ell=2$ and even function $F_{0}(r)$, the solution is
$F(r,t)$ $=$ $\frac{1}{2r}\left\{(r-t)F_{0}(r-t)+(r+t)F_{0}(r+t)\right\}$
$+\frac{1}{2r^{3}}t\Bigg{\{}-\frac{3}{2}(r^{2}-t^{2})\int_{r-t}^{r+t}\frac{F_{0}(\tau)}{\tau}\,d\tau-\frac{3}{2}\int_{r-t}^{r+t}\tau F_{0}(\tau)\,d\tau\Bigg{\}}\,.$
For $\ell=3$ and odd function $F_{0}(r)$, the solution is
$F(r,t)$ $=$ $\frac{1}{2r}\left\{(r-t)F_{0}(r-t)+(r+t)F_{0}(r+t)\right\}$
$\frac{1}{2r^{4}}t\left\{-\frac{15}{8}(r^{2}-t^{2})^{2}\int_{r-t}^{r+t}\frac{F_{0}(\tau)}{\tau^{2}}d\tau\right$.
$\left.-\frac{1}{4}\left(9r^{2}-15t^{2}\right)\int_{r-t}^{r+t}F_{0}(\tau)d\tau-\frac{15}{8}\int_{r-t}^{r+t}\tau^{2}F_{0}(\tau)d\tau\right\}\,.$
For $\ell=4$ and even function $F_{0}(r)$, the solution is
$F(r,t)$ $=$ $\frac{1}{2r}\left\{(r-t)F_{0}(r-t)+(r+t)F_{0}(r+t)\right\}$
$-\frac{1}{2r^{5}}t\left\{\frac{35}{16}\left(r^{6}-3r^{4}t^{2}+3r^{2}t^{4}-t^{6}\right)\int_{r-t}^{r+t}\frac{F_{0}(\tau)}{\tau^{3}}d\tau\right$.
$+\frac{5}{16}\left(9r^{4}-30r^{2}t^{2}+21t^{4}\right)\int_{r-t}^{r+t}\frac{F_{0}(\tau)}{\tau}d\tau$
$\left.+\frac{5}{16}\left(9r^{2}-21t^{2}\right)\int_{r-t}^{r+t}\tau F_{0}(\tau)d\tau+\frac{35}{16}\int_{r-t}^{r+t}\tau^{3}F_{0}(\tau)d\tau\right\}\,.$
The obtained formulas allow us to estimate the “radial tail” of the solution.
Lemma 2.1
Assume that $F_{1}(r)\equiv 0$. For a pionic atom with (2.5), the tail undergoes an exponential decay in time
$\left|F(r,t)-\frac{1}{2r}\left\{(r-t)F_{0}(r-t)+(r+t)F_{0}(r+t)\right\}\right|\leq C(r)e^{-t/2}t^{n+\mu-\frac{3}{2}},\quad\ell=1,2,\ldots\,.$
Proof. Indeed, for even $\ell$ the tail contains the integrals
$\int_{r-t}^{r+t}F_{0}(\tau)\tau^{k}d\tau$ $=$ $\int_{t-r}^{t+r}F_{0}(\tau)\tau^{k}d\tau,\quad k=-\ell+1,-\ell+3,\ldots,\ell-3,\ell-1\,.$
For a pionic atom with (2.5) for every given $r$ and a sufficiently large time $t>r+1$,
$\left|\int_{r-t}^{r+t}F_{0}(\tau)\tau^{k}d\tau\right|$ $\leq$ $\int_{t-r}^{t+r}\left|F_{0}(\tau)\tau^{k}\right|d\tau$
$\leq$ $C_{n,\ell,\mu}\int_{t-r}^{t+r}\left|\tau^{\mu-\frac{1}{2}}e^{-\tau/2}L_{n-\ell-1}^{2\mu}(\tau)\tau^{k}\right|d\tau$
$\leq$ $C_{n,\ell,\mu}\int_{t-r}^{t+r}e^{-\tau/2}\tau^{a}d\tau\,$,
where $a:=\mu-\frac{1}{2}+n-\ell-1+k$. The last integral can be evaluated in terms of the incomplete gamma function ([3, Ch.9]):
$\left|\int_{r-t}^{r+t}F_{0}(\tau)\tau^{k}d\tau\right|$ $\leq$ $C_{n,\ell,\mu}2^{a+1}\left|\Gamma\left(a+1,\frac{t-r}{2}\right)-\Gamma\left(a+1,\frac{r+t}{2}\right)\right|$
$\leq$ $2C_{n,\ell,\mu}\Bigg{|}\left(1+O\left(\frac{1}{t}\right)\right)(t-r)^{a}\exp\left(-\frac{t}{2}+\frac{r}{2}+O\left(\frac{1}{t^{2}}\right)\right)$
$-\left(1+O\left(\frac{1}{t}\right)\right)(r+t)^{a}\exp\left(-\frac{t}{2}-\frac{r}{2}+O\left(\frac{1}{t^{2}}\right)\right)\Bigg{|}\,.$
Hence,
$\left|F(r,t)-\frac{1}{2r}\left\{(r-t)F_{0}(r-t)+(r+t)F_{0}(r+t)\right\}\right|$ $\leq$ $C(r)te^{-t/2}\sum_{k\,odd,\,\,k=-\ell+1}^{\ell-1}t^{\ell-1-k}t^{a}$
$\leq$ $C(r)e^{-t/2}t^{n+\mu-\frac{3}{2}}\,.$
For odd $\ell$, the arguments are similar. The lemma is proved. $\square$
The solutions with nontrivial $F_{1}$ and their estimate can be easily derived from these formulas by integration over time.
Although the equation (2.3) received the treatment in the textbooks, its application to wave propagation in cosmological backgrounds can still be extended. One must refer to [28], where the author considered the problem of gravitational radiation emitted by a small particle falling into the Schwarzschild field and examined its spectrum in the high-frequency limit. In [17], the author reviewed equation (2.3) and applied it to the computation of the leading terms in the far-field backscatter from outgoing radiation in Schwarzschild spacetime and of the matter-induced singularities in plane symmetric spacetimes.
2.2 The Riemann function approach
The recursive procedures (2.6),(2.1), and (2.8), can be reduced to the single formula if we appeal to the Riemann function and modify conditions on the functions at infinity. Indeed, the equation (2.3) can be written in the form that leads to the problem for the unknown function $R(r,t)=rF(r,t)$. In this subsection, we provide an explicit formula for the solution to the Cauchy problem
$\cases{\displaystyle R_{tt}(r,t)-R_{rr}(r,t)+\frac{\ell(\ell+1)}{r^{2}}R(r,t)=0,\quad r>0,\quad t\geq 0\cr{}\cr R(r,0)=R_{0}(r),\quad R_{t}(r,0)=R_{1}(r),\quad r>0,\cr{}\cr R(0,t)=0,\quad\quad t\geq 0.}$ (2.9)
For the Klein-Gordon equation, we set the initial functions as follows:
$R_{k}(r)=O(r^{\mu+\frac{1}{2}})\,\,\mbox{\rm as }\,\,r\searrow 0,\quad k=0,1\,.$ (2.10)
The functions $R_{0}(r)$ and $R_{1}(r)$ will be continued to negative $r$ as odd functions for even $\ell$ and as even functions for odd $\ell$.
Lemma 2.2
Let $\ell$ be positive integer number and assume that $R_{0}=R_{0}(r)$ and $R_{1}=R_{1}(r)$ are odd functions for even $\ell$ and they are even functions for odd $\ell$. Assume also $R_{0},R_{1}\in C^{\infty}({\mathbb{R}}_{+})$, condition (2.10) is fulfilled and that $\mu>\ell-\frac{3}{2}$. Then the solution to the problem (2.9) with the property (2.4) is given by
$R(r,t)$
$=$ $\!\!\frac{1}{2}[R_{0}(r-t)+R_{0}(r+t)]-\frac{1}{4}\ell(\ell+1)\frac{t}{r}\int_{r-t}^{r+t}R_{0}(s)\frac{1}{s}F\left(1-\ell,\ell+2;2;\frac{t^{2}-(r-s)^{2}}{4rs}\right)ds$
$+\int_{0}^{t}\frac{1}{2}[R_{1}(r-\tau)+R_{1}(r+\tau)]\,d\tau$
$-\int_{0}^{t}\frac{1}{4}\ell(\ell+1)\frac{\tau}{r}\int_{r-\tau}^{r+\tau}R_{1}(s)\frac{1}{s}F\left(1-\ell,\ell+2;2;\frac{\tau^{2}-(r-s)^{2}}{4rs}\right)\,ds\,d\tau.$
Proof. We skip the proof since it is standard and can be easily reconstructed. (See, e.g., [17, 22] and references therein.) $\square$
In particular, for $\ell=5$, in addition to the cases obtained in subsection 2.1, we obtain
$F(r,t)$
$=$ $\frac{1}{2r}[(r-t)F_{0}(r-t)+(r+t)F_{0}(r+t)]$
$+\frac{1}{2r^{6}}t\left\{\frac{1}{2^{7}}\left(-315r^{8}+1260r^{6}t^{2}-1890r^{4}t^{4}+1260r^{2}t^{6}-315t^{8}\right)\int_{r-t}^{r+t}F_{0}(\tau)\frac{1}{\tau^{4}}\,d\tau\right$.
$+\frac{1}{2^{5}}\left(-105r^{6}+525r^{4}t^{2}-735r^{2}t^{4}+315t^{6}\right)\int_{r-t}^{r+t}F_{0}(\tau)\frac{1}{\tau^{2}}\,d\tau$
$+\frac{1}{2^{6}}\left(-225r^{4}+1050r^{2}t^{2}-945t^{4}\right)\int_{r-t}^{r+t}F_{0}(\tau)\,d\tau$
$\left.+\frac{1}{2^{5}}\left(315t^{2}-105r^{2}\right)\int_{r-t}^{r+t}F_{0}(\tau)\tau^{2}\,d\tau-\frac{315}{2^{7}}\int_{r-t}^{r+t}F_{0}(\tau)\tau^{4}\,d\tau\right\}\,.$
Remark 2.3
The formula of Lemma 2.2 as well as the recursive procedures (2.6), (2.1), and (2.8) exhibit the finite propagation speed property. In particular, for a fixed time, in the fixed bounded domain, the values at infinity of functions $R=R_{0}(r)$ and $R=R_{1}(r)$ do not affect the solution.
2.3 Solution by Hankel transform
[…]
$F_{\nu}(s):=\int_{0}^{\infty}rf(r)J_{\nu}(sr)\,dr\,.$
with $\nu>-1/2$ ([16, Sec.11.5.1], [19, Sec.14.3], [27, Sec.14.3]). Let $f(r)$ be a function defined for $r>0$. The inversion formula yields
$f(r)=\int_{0}^{\infty}sF_{\nu}(s)J_{\nu}(sr)\,ds\,.$
Let $Y_{\ell m}(\theta,\phi)$, $|m|\leq\ell$, $\ell=0,1,\ldots$, be spherical harmonics. We skip the proof of the following statement.
Lemma 2.4
Assume that $F_{0}(r),F_{1}(r)\in C^{1}(0,\infty)$, and
$\int_{0}^{\infty}r^{3/2}|F_{i}(r)|\,dr<\infty,\quad F_{i}(r)=O(r^{-k}),\quad r\to\infty,\quad k>2,\quad i=0,1.$
Then the solution to the problem
$V_{tt}-\bigtriangleup V=0,\quad V(x,0)=0,\quad V_{t}(x,0)=F_{1}(r)Y_{\ell m}(\theta,\phi)\,,\quad t\in{\mathbb{R}}_{+},\,x\in{\mathbb{R}}^{3}\,$,
is
$V(r,\theta,\phi,t)$ $=$ $Y_{\ell m}(\theta,\phi)\frac{1}{\sqrt{r}}\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{1}(\rho)J_{\ell+\frac{1}{2}}(\lambda\rho)\,\rho^{3/2}d\rho\right)\sin(\lambda t)J_{\ell+\frac{1}{2}}(r\lambda)\,d\lambda\,$,
while the solution to the problem
$v_{tt}-\bigtriangleup v=0,\quad v(x,0)=F_{0}(r)Y_{\ell m}(\theta,\phi),\quad v_{t}(x,0)=0\,$,
is given by
$v(r,\theta,\phi,t)$ $=$ $Y_{\ell m}(\theta,\phi)\frac{1}{\sqrt{r}}\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{0}(\rho)J_{\ell+\frac{1}{2}}(\rho\lambda)\rho^{3/2}\,d\rho\right)\cos\left(\lambda t\right)J_{\ell+\frac{1}{2}}(r\lambda)\lambda\,d\lambda\,.$
In particular, we can set
$R(-r,t)$ $:=$ $(-1)^{\ell+1}R(r,t)\quad\mbox{\rm for all}\quad r>0\,.$
Thus, the kernel $K(r,t;\rho)$ of the integral transform
$R(r,t)$ $=$ $\int_{0}^{\infty}R_{1}(\rho)K(r,t;\rho)\,\,d\rho,\quad r>0,\quad t\geq 0\,$,
is
$K(r,t;\rho)$ $=$ $\int_{0}^{\infty}\sin(\lambda t)\sqrt{\rho}J_{\ell+\frac{1}{2}}(\lambda\rho)\sqrt{r}J_{\ell+\frac{1}{2}}(r\lambda)\,\,d\lambda\,,\quad r>0\,,\quad\rho>0\,,\quad t\geq 0\,.$
Note that the function
$k(r,t;\rho,\lambda)$ $=$ $\sin(\lambda t)\sqrt{\rho}J_{\ell+\frac{1}{2}}(\lambda\rho)\sqrt{r}J_{\ell+\frac{1}{2}}(r\lambda)\,,\quad r>0,\,\,t>0,\,\,\lambda>0,\,\,\rho>0\,$,
solves the equation
$\displaystyle k_{tt}(r,t;\rho,\lambda)-k_{rr}(r,t;\rho,\lambda)+\frac{\ell(\ell+1)}{r^{2}}k(r,t;\rho,\lambda)=0,\quad t>0,\quad r>0\,.$
The isometry property in the weighted Sobolev spaces of the Hankel transform is studied in [1, 20], and it can be extended to the problem of Theorem 1.2.
3 Proof of Theorems 1.1,1.2
To illustrate ITA, we first present the case of the Minkowski space. Consider in the Minkowski space the spherical Klein-Gordon fields, which, for instance, can be emitted by the wave function of the pionic atom. Application of the ITA to the Klein-Gordon equation in the Minkowski space leads to the following formula
$u(x,t)$ $=$ $v_{\varphi_{0}}(x,t)$
$-\int_{0}^{t}\frac{m_{0}t}{\sqrt{t^{2}-\tau^{2}}}J_{1}\left(m_{0}\sqrt{t^{2}-\tau^{2}}\right)v_{\varphi_{0}}(x,\tau)\,d\tau+\int_{0}^{t}J_{0}\left(m_{0}\sqrt{t^{2}-\tau^{2}}\right)v_{\varphi_{1}}(x,\tau)\,d\tau$
for the solution of the problem
$u_{tt}-\Delta u+m_{0}^{2}u=0\quad\mbox{\rm in}\quad{\mathbb{R}}^{n+1},\quad u(x,0)=\varphi_{0}(x),\quad u_{t}(x,0)=\varphi_{1}(x)\quad\mbox{\rm in}\quad{\mathbb{R}}^{n}\,.$
Here $v_{\varphi}(x,r)$ is a solution to the problem
$v_{tt}-\Delta v=0\quad\mbox{\rm in}\quad{\mathbb{R}}^{n+1},\quad v(x,0)=\varphi(x),\quad v_{t}(x,0)=0\quad\mbox{\rm in}\quad{\mathbb{R}}^{n}\,.$
One can use any form of the spherical waves given in Section 2. We choose Lemma 2.4 and skip an elementary proof of the following statement.
Theorem 3.1
Assume that $F_{0}(r),F_{1}(r)\in C^{1}(0,\infty)$, and
$\int_{0}^{\infty}r^{3/2}|F_{i}(r)|\,dr<\infty,\quad F_{i}(r)=O(r^{-k}),\quad r\to\infty,\quad k>2,\quad i=0,1.$
The spherical Klein-Gordon field in the Minkowski space is given by
$\Phi(r,\theta,\phi,t)$ $=$ $Y_{\ell m}(\theta,\phi)\frac{1}{\sqrt{r}}\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{0}(\rho)J_{\ell+\frac{1}{2}}(\rho\lambda)\rho^{3/2}\,d\rho\right)\cos\left(\lambda t\right)J_{\ell+\frac{1}{2}}(r\lambda)\lambda\,d\lambda$
$-Y_{\ell m}(\theta,\phi)\frac{1}{\sqrt{r}}\int_{0}^{t}\Bigg{\{}\frac{m_{0}t}{\sqrt{t^{2}-\tau^{2}}}J_{1}\left(m_{0}\sqrt{t^{2}-\tau^{2}}\right)-J_{0}\left(m_{0}\sqrt{t^{2}-\tau^{2}}\right)\Bigg{\}}$
$\times\Bigg{\{}\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{1}(\rho)J_{\ell+\frac{1}{2}}(\rho\lambda)\rho^{3/2}\,d\rho\right)\cos\left(\lambda\tau\right)J_{\ell+\frac{1}{2}}(r\lambda)\lambda\,d\lambda\Bigg{\}}\,d\tau$
for all $x\in{\mathbb{R}}^{3}$, $t>0$.
Next we turn to the de Sitter universe. The ITA applied to the initial value problem (1.2) (written in the Cartesian coordinates $x\in{\mathbb{R}}^{n}$), for the equation
$\frac{\partial^{2}\psi}{\partial t^{2}}+nH\frac{\partial\psi}{\partial t}-e^{-2Ht}{\mathcal{A}}(x,\partial_{x})\psi+\frac{m_{n}^{2}}{\hbar^{2}}\psi=0\,$, (3.1)
with the operator ${\mathcal{A}}(x,\partial_{x})=\sum_{|\alpha|\leq 2}a_{\alpha}(x)\partial_{x}^{\alpha}$, leads to the formula for the solution
$\psi(x,t)$
$=$ $e^{-\frac{n-1}{2}Ht}v_{\psi_{0}}(x,\phi(t))+\,e^{-\frac{n}{2}Ht}\int_{0}^{\phi(t)}\left[2K_{0}(s,t;M)+nHK_{1}(s,t;M)\right]v_{\psi_{0}}(x,s)ds$
+ $+\,2e^{-\frac{n}{2}Ht}\int_{0}^{\phi(t)}v_{\psi_{1}}(x,s)K_{1}(s,t;M)ds,\quad\forall x\in\Omega\subseteq{\mathbb{R}}^{n}\,,\,\,\forall t\in I=[0,T]\subseteq[0,\infty),$
of the Cauchy problem for the equation (3.1) with the initial data $\psi(x,0)=\psi_{0}(x)$ and $\psi_{t}(x,0)=\psi_{1}(x)$. Here the kernels $K_{0}(s,t;M)$ and $K_{1}(s,t;M)$, are defined in (1) and (1), (see [22, 26]), respectively, where $0<T\leq\infty$, $\phi(t):=(1-e^{-Ht})/H$, $M^{2}=\frac{n^{2}H^{2}}{4}-\frac{m^{2}c^{4}}{\hbar^{2}}\in{\mathbb{C}}$, and the function   $v=v_{\varphi}(x,t)$   is a solution of the problem
$\cases{v_{tt}-{\mathcal{A}}(x,\partial_{x})v=0,\quad\forall x\in\Omega\,,\quad\forall t\in[0,(1-e^{-HT})/H]\,,\cr v(x,0)=\varphi(x),\quad v_{t}(x,0)=0\,,\quad\forall x\in\Omega\,.}$ (3.3)
Here, the function $v_{\varphi}(x,\phi(t))$ coincides with the value $v(x,\phi(t))$ of the solution $v(x,t)$ of the Cauchy problem (3.3) with ${\mathcal{A}}(x,\partial_{x})=\Delta$. In the case of $n=3$ and $m_{n}^{2}=2H^{2}\hbar^{2}$ the field is huygensian, that is, it obeys the Huygens’ principle (see [23]) and the kernels are simplified as follows:
$K_{0}\left(r,t;\frac{1}{2}H\right)=-\frac{1}{4}He^{\frac{Ht}{2}}\,,\quad K_{1}\left(r,t;\frac{1}{2}H\right)=\frac{1}{2}e^{\frac{Ht}{2}}\,.$
(For the definition of the Huygens’ principle see, e.g., [11].)
3.1 Proof of Theorem 1.1
From (3) we derive
$\Phi(r,\theta,\phi,t)$ $=$ $e^{-Ht}v_{F_{0}Y_{\ell m}}\left(x,\phi(t)\right)+\,e^{-H\frac{3}{2}t}\int_{0}^{\phi(t)}v_{F_{0}Y_{\ell m}}\left(x,\tau\right)\left(2K_{0}\left(\tau,t\right)+3K_{1}\left(\tau,t\right)\right)\,d\tau$ (3.4)
+ $+\,2e^{-\frac{3}{2}Ht}\int_{0}^{\phi(t)}v_{F_{1}Y_{\ell m}}\left(x,\tau\right)K_{1}\left(\tau,t\right)\,d\tau\,.$
Then we use the formula from Lemma 2.2 with $F_{1}=0$, that is,
$v_{F_{k}Y_{\ell m}}\left(x,t\right)$ $=$ $Y_{\ell m}(\theta,\phi)\Bigg{\{}\frac{1}{2r}[(r-t)F_{k}(r-t)+(r+t)F_{k}(r+t)]$
$-\frac{1}{4}\ell(\ell+1)\frac{t}{r^{2}}\int_{r-t}^{r+t}F_{k}(s)F\left(1-\ell,\ell+2;2;\frac{t^{2}-(r-s)^{2}}{4rs}\right)ds\Bigg{\}},\quad k=0,1\,.$
Hence, with $\phi(t):=\frac{1-e^{-Ht}}{H}$ and $\tau:=\phi(t)s=\frac{1-e^{-Ht}}{H}s$ we obtain
$\Phi(r,\theta,\phi,t)$
$=$ $e^{-Ht}Y_{\ell m}(\theta,\phi)\Bigg{\{}\frac{1}{2r}[(r-\phi(t))F_{0}(r-\phi(t))+(r+\phi(t))F_{0}(r+\phi(t))]$
$-\frac{1}{4}\ell(\ell+1)\frac{\phi(t)}{r^{2}}\int_{r-\phi(t)}^{r+\phi(t)}F_{0}(s)F\left(1-\ell,\ell+2;2;\frac{\phi(t)^{2}-(r-s)^{2}}{4rs}\right)ds\Bigg{\}}$
$+\,e^{-H\frac{3}{2}t}Y_{\ell m}(\theta,\phi)\int_{0}^{\phi(t)}\Bigg{\{}\frac{1}{2r}[(r-\tau)F_{0}(r-\tau)+(r+\tau)F_{0}(r+\tau)]$
$-\frac{1}{4}\ell(\ell+1)\frac{\tau}{r^{2}}\int_{r-\tau}^{r+\tau}F_{0}(s)F\left(1-\ell,\ell+2;2;\frac{t^{2}-(r-s)^{2}}{4rs}\right)ds\Bigg{\}}\left(2K_{0}\left(\tau,t\right)+3K_{1}\left(\tau,t\right)\right)\,d\tau$
+ $+\,2e^{-\frac{3}{2}Ht}Y_{\ell m}(\theta,\phi)\int_{0}^{\phi(t)}\Bigg{\{}\frac{1}{2r}[(r-\tau)F_{1}(r-\tau)+(r+\tau)F_{1}(r+\tau)]$
$-\frac{1}{4}\ell(\ell+1)\frac{\tau}{r^{2}}\int_{r-\tau}^{r+\tau}F_{1}(s)F\left(1-\ell,\ell+2;2;\frac{\tau^{2}-(r-s)^{2}}{4rs}\right)ds\Bigg{\}}K_{1}\left(\tau,t\right)\,d\tau\,.$
For the huygensian field with $\varphi_{0}(x)=\Phi(x,0)$ and $\varphi_{1}(x)=\Phi_{t}(x,0)$ of (1.2), the field in the de Sitter FLRW spacetime is (see [23])
$\Phi(r,\theta,\phi,t)$ $=$ $e^{-Ht}v_{\varphi_{0}}\left(x,\phi(t)\right)+\,e^{-Ht}V_{\varphi_{0}+\varphi_{1}}\left(x,\phi(t)\right)\,$, (3.6)
where $V_{\varphi_{0}+\varphi_{1}}(x,t)$ is a solution of the problem
$V_{tt}-\Delta V=0,\quad V(x,0)=0,\quad V_{t}(x,0)=\varphi_{0}(x)+\varphi_{1}(x)\,$,
and $V_{\varphi_{0}+\varphi_{1}}(x,t)=\int_{0}^{t}v(x,\tau)\,d\tau$. Next, we use (3.1),
$v_{F_{0}Y_{\ell m}}\left(x,\phi(t)\right)$ $=$ $Y_{\ell m}(\theta,\phi)\Bigg{\{}\frac{1}{2r}[(r-\phi(t))F_{0}(r-\phi(t))+(r+\phi(t))F_{0}(r+\phi(t))]$
$-\frac{1}{4}\ell(\ell+1)\frac{\phi(t)}{r^{2}}\int_{r-\phi(t)}^{r+\phi(t)}F_{0}(s)F\left(1-\ell,\ell+2;2;\frac{\phi(t)^{2}-(r-s)^{2}}{4rs}\right)ds\Bigg{\}},$
and
$V_{(F_{0}+F_{1})Y_{\ell m}}\left(x,\phi(t)\right)$
$=$ $Y_{\ell m}(\theta,\phi)\int_{0}^{\phi(t)}\Bigg{\{}\frac{1}{2r}\Big{(}(r-\tau)[F_{0}(r-\tau)+F_{1}(r-\tau)]+(r+\tau)[F_{0}(r+\tau)+F_{1}(r+\tau)]\Big{)}$
$-\frac{1}{4}\ell(\ell+1)\frac{\tau}{r^{2}}\int_{r-\tau}^{r+\tau}(F_{0}(s)+F_{1}(s))F\left(1-\ell,\ell+2;2;\frac{\tau^{2}-(r-s)^{2}}{4rs}\right)ds\Bigg{\}}d\tau,$
to obtain the last statement of the theorem. The theorem is proved. $\square$
3.2 Proof of Theorem 1.2
Proof of Theorem 1.2. If we use the formula (3.4), then we derive
$\Phi(r,\theta,\phi,t)$
$=$ $e^{-Ht}Y_{\ell m}(\theta,\phi)\frac{1}{\sqrt{r}}\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{0}(\rho)J_{\ell+\frac{1}{2}}(\rho\lambda)\rho^{3/2}\,d\rho\right)\cos\left(\lambda\phi(t)\right)J_{\ell+\frac{1}{2}}(r\lambda)\lambda\,d\lambda$
$+\,e^{-H\frac{3}{2}t}Y_{\ell m}(\theta,\phi)\frac{1}{\sqrt{r}}\int_{0}^{1}\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{0}(\rho)J_{\ell+\frac{1}{2}}(\rho\lambda)\rho^{3/2}\,d\rho\right)\cos\left(\lambda\phi(t)s\right)$
$\times J_{\ell+\frac{1}{2}}(r\lambda)\lambda\,d\lambda\left(2K_{0}\left(\phi(t)s,t\right)+3K_{1}\left(\phi(t)s,t\right)\right)\phi(t)\,ds$
+ $+\,2e^{-\frac{3}{2}Ht}Y_{\ell m}(\theta,\phi)\frac{1}{\sqrt{r}}\int_{0}^{1}\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{1}(\rho)J_{\ell+\frac{1}{2}}(\rho\lambda)\rho^{3/2}\,d\rho\right)\cos\left(\lambda\phi(t)s\right)$
$\times J_{\ell+\frac{1}{2}}(r\lambda)\lambda\,d\lambda\,K_{1}\left(\phi(t)s,t\right)\phi(t)\,ds\quad\mbox{\rm for all}\quad t>0\,.$
For the huygensian field we use (3.6) and the formula from Lemma 2.4, to obtain
$\Phi(r,\theta,\phi,t)$
$=$ $e^{-Ht}Y_{\ell m}(\theta,\phi)\frac{1}{\sqrt{r}}\Bigg{\{}\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{0}(\rho)J_{\ell+\frac{1}{2}}(\rho\lambda)\rho^{3/2}\,d\rho\right)\cos\left(\lambda\phi(t)\right)J_{\ell+\frac{1}{2}}(r\lambda)\lambda\,d\lambda$
$+\int_{0}^{\infty}\left(\int_{0}^{\infty}(F_{0}(\rho)+F_{1}(\rho))J_{\ell+\frac{1}{2}}(\lambda\rho)\,\rho^{3/2}d\rho\right)\sin\left(\lambda\phi(t)\right)J_{\ell+\frac{1}{2}}(r\lambda)\,d\lambda\Bigg{\}}\,.$
Theorem 1.2 is proved. $\square$
Corollary 3.2
Let $F_{1}(r)\equiv 0$ and $F_{0}(r)$ be of the form (2.5). If $3H\hbar/2\leq m_{n}$, then
$\left|\Phi(r,\theta,\phi,t)-e^{-Ht}v_{F_{0}Y_{\ell m}}(r,\theta,\phi,\phi(t))\right|$ $\leq$ $C(r)e^{-Ht}\left|Y_{\ell m}(\theta,\phi)\right|(1+t)^{1-\mathop{\mathrm{sgn}}(|M|)}\,$,
where $\phi(t)=(1-e^{-Ht})/H$. If $3H\hbar/2>m_{n}$, then
$\left|\Phi(r,\theta,\phi,t)-e^{-Ht}v_{F_{0}Y_{\ell m}}(r,\theta,\phi,\phi(t))\right|$
$\leq$ $C(r)e^{-\frac{3}{2}Ht}\left|Y_{\ell m}(\theta,\phi)\right|\cases{e^{\frac{1}{2}Ht}\quad\mbox{\rm if}\quad 3H\hbar/2>m_{n}>\sqrt{2}H\hbar\,,\cr e^{t\sqrt{\frac{9}{4}H^{2}-\frac{m_{n}^{2}}{\hbar^{2}}}}\quad\mbox{\rm if}\quad m_{n}<\sqrt{2}H\hbar\,.}$
If $m_{n}=\sqrt{2}H\hbar$, then
$\left|\psi(x,t)-e^{-Ht}v_{\psi_{0}}(x,\phi(t))\right|$ $\leq$ $C(r)e^{-Ht}\left|Y_{\ell m}(\theta,\phi)\right|\,.$
Here $v_{F_{0}Y_{\ell m}}(x,\phi(t))$ is given by Lemma 2.4,
$v_{F_{0}Y_{\ell m}}(r,\theta,\phi,\phi(t))$
$=$ $Y_{\ell m}(\theta,\phi)\frac{1}{\sqrt{r}}\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{0}(\rho)J_{\ell+\frac{1}{2}}(\rho\lambda)\rho^{3/2}\,d\rho\right)\cos\left(\lambda\phi(t)\right)J_{\ell+\frac{1}{2}}(r\lambda)\lambda\,d\lambda\,.$
Proof. We appeal to the representation (3) with the dimension of spatial variable $n=3$,
$\psi(x,t)=e^{-Ht}v_{\psi_{0}}(x,\phi(t))+\,e^{-\frac{3}{2}Ht}\int_{0}^{\phi(t)}\left[2K_{0}(s,t;M)+3HK_{1}(s,t;M)\right]v_{\psi_{0}}(x,s)ds\,$,
and to Lemma 2.1. For the tail, this implies
$\left|\Phi(r,\theta,\phi,t)-e^{-Ht}v_{F_{0}Y_{\ell m}}(x,\phi(t))\right|$
$=$ $\left|e^{-\frac{3}{2}Ht}\int_{0}^{\phi(t)}\left[2K_{0}(s,t;M)+3HK_{1}(s,t;M)\right]v_{F_{0}Y_{\ell m}}(x,s)ds\right|$
$\leq$ $e^{-\frac{3}{2}Ht}\left|Y_{\ell m}(\theta,\phi)\right|\int_{0}^{\phi(t)}\left|2K_{0}(s,t;M)+3HK_{1}(s,t;M)\right|$
$\times\left(\frac{1}{2r}\left|(r-s)F_{0}(r-s)+(r+s)F_{0}(r+s)\right|+C(r)e^{-s/2}s^{n+\mu-\frac{3}{2}}\right)ds$
$\leq$ $C(r)e^{-\frac{3}{2}Ht}\left|Y_{\ell m}(\theta,\phi)\right|\int_{0}^{\phi(t)}\left|2K_{0}(s,t;M)+3HK_{1}(s,t;M)\right|$
$\times\Bigg{(}\frac{1}{2r}\Bigg{[}|r-s|^{\mu+\frac{1}{2}}e^{-|r-s|/2}|L_{n-\ell-1}^{2\mu}(|r-s|)|$
$+|r+s|^{\mu+\frac{1}{2}}e^{-|r+s|/2}|L_{n-\ell-1}^{2\mu}(|r+s|)|\Bigg{]}+e^{-s/2}s^{n+\mu-\frac{3}{2}}\Bigg{)}ds$
$\leq$ $C(r)e^{-\frac{3}{2}Ht}\left|Y_{\ell m}(\theta,\phi)\right|\int_{0}^{\phi(t)}\left|2K_{0}(s,t;M)+3HK_{1}(s,t;M)\right|$
$\times\Bigg{(}(r+s)^{\mu+\frac{1}{2}+n-\ell-1}+e^{-s/2}s^{n+\mu-\frac{3}{2}}\Bigg{)}ds$
$\leq$ $C(r)e^{-\frac{3}{2}Ht}\left|Y_{\ell m}(\theta,\phi)\right|\int_{0}^{\phi(t)}\left|2K_{0}(s,t;M)+3HK_{1}(s,t;M)\right|ds,\quad\ell=1,2,\ldots,$
where $v_{F_{0}Y_{\ell m}}(x,\phi(t))$ can be written in three different explicit forms of Subsections 2.1-2.3. Recall, $M:=\left(\frac{9}{4}H^{2}-\frac{m_{n}^{2}}{\hbar^{2}}\right)^{1/2}$; hence, either $M\geq 0$ or $\Re M=0$. According to Lemmas 4.2, 4.3 [26], if $M\geq 0$ and $H$ is a positive number, then
$\int_{0}^{\phi(t)}|K_{0}(s,t;-iM)|ds$ $\leq$ $C_{M,H}(1+t)^{1-\mathop{\mathrm{sgn}}(M)}e^{-Ht/2}(e^{Ht}-1)\quad\mbox{for all}\quad t\in[0,\infty)\,$,
$\int_{0}^{\phi(t)}|K_{1}(r,t;-iM)|dr$ $\leq$ $C_{M}(1+t)^{1-\mathop{\mathrm{sgn}}(M)}\frac{1}{H}(e^{Ht}-1)(e^{Ht}+1)^{-1}\quad\mbox{for all}\quad t\in[0,\infty)\,.$
According to Lemmas 5.1, 5.2 [26], if a>1a>-1, $M>0$, and $\phi(t)=(1-e^{-Ht})/H$, then for large $t$,
$\int_{0}^{1}\phi(t)^{a}s^{a}\big{|}K_{1}(\phi(t)s,t;M)\big{|}\phi(t)\,ds$ $\leq$ $C_{M,a}e^{Mt}\,$,
$\int_{0}^{1}\phi(t)^{a}s^{a}\big{|}K_{0}(\phi(t)s,t;M)\big{|}\phi(t)\,ds$ $\leq$ $C_{M,a}\times\cases{e^{\frac{1}{2}Ht}\quad\mbox{\rm if}\quad M<H/2\,,\cr e^{(M-H)t}\quad\mbox{\rm if}\quad M>H/2\,.}$
Consequently, if $3H\hbar/2>m_{n}$, then
$\left|\Phi(r,\theta,\phi,t)-e^{-Ht}v_{F_{0}Y_{\ell m}}(x,\phi(t))\right|$
$\leq$ $C(r)e^{-\frac{3}{2}Ht}\left|Y_{\ell m}(\theta,\phi)\right|$
$\times\left(e^{\left(\frac{9}{4}H^{2}-\frac{m_{n}^{2}}{\hbar^{2}}\right)^{1/2}t}+\cases{e^{\frac{1}{2}Ht}\quad\mbox{\rm if}\quad 3H\hbar/2>m_{n}>\sqrt{2}H\hbar\,,\cr e^{(\left(\frac{9}{4}H^{2}-\frac{m_{n}^{2}}{\hbar^{2}}\right)^{1/2}-H)t}\quad\mbox{\rm if}\quad m_{n}<\sqrt{2}H\hbar\,,}\right)$
$\leq$ $C(r)e^{-\frac{3}{2}Ht}\left|Y_{\ell m}(\theta,\phi)\right|\cases{e^{\frac{1}{2}Ht}\quad\mbox{\rm if}\quad 3H\hbar/2>m_{n}>\sqrt{2}H\hbar\,,\cr e^{t\sqrt{\frac{9}{4}H^{2}-\frac{m_{n}^{2}}{\hbar^{2}}}}\quad\mbox{\rm if}\quad m_{n}<\sqrt{2}H\hbar\,.}$
If $3H\hbar/2\leq m_{n}$, then
$\left|\Phi(r,\theta,\phi,t)-e^{-Ht}v_{F_{0}Y_{\ell m}}(x,\phi(t))\right|$
$\leq$ $C(r)e^{-\frac{3}{2}Ht}\left|Y_{\ell m}(\theta,\phi)\right|\int_{0}^{\phi(t)}\left|2K_{0}(s,t;M)+3HK_{1}(s,t;M)\right|ds$
$\leq$ $C(r)e^{-Ht}\left|Y_{\ell m}(\theta,\phi)\right|(1+t)^{1-\mathop{\mathrm{sgn}}(|M|)}\,.$
For the case of huygensian field $m_{n}=\sqrt{2}H\hbar$, the solution (3.6) implies
$\left|\Phi(r,\theta,\phi,t)-e^{-Ht}v_{\psi_{0}}(x,\phi(t))\right|$
$\leq$ $e^{-Ht}H\int_{0}^{\phi(t)}\left|v_{\psi_{0}}(x,s)ds\right|$
$\leq$ $e^{-Ht}H\left|Y_{\ell m}(\theta,\phi)\right|$
$\times\int_{0}^{\phi(t)}\left|\frac{1}{2r}\left|(r-s)F_{0}(r-s)+(r+s)F_{0}(r+s)\right|+C(r)e^{-s/2}s^{n+\mu-\frac{3}{2}}\right|ds$
$\leq$ $C(r)e^{-Ht}\left|Y_{\ell m}(\theta,\phi)\right|\,.$
The corollary is proved. $\square$
For the pionic atom, the modification is evident since $F_{1}(r)=-\frac{i}{\hbar}E_{n}F_{0}(r)$. If $\ell=0$, then the explicit formula of Subsection 2.1 can be used to estimate a solution. In particular, for the ground state $n=1$ and $\ell=0$, in the case of huygensian field $m_{n}=\sqrt{2}H\hbar$, according to (3.6), at $r=1/H$, we obtain
$\Phi\left(\frac{1}{H},\theta,\phi,t\right)$ $=$ $e^{-Ht}H^{\frac{1}{2}-\mu}\frac{1}{4\sqrt{\pi}}\Bigg{[}\Bigg{\{}e^{-H(\mu+\frac{1}{2})t-\frac{1}{2H}e^{-Ht}}+\left(2-e^{-Ht}\right)^{\mu+\frac{1}{2}}e^{-\frac{1}{2H}(2-e^{-Ht})}\Bigg{\}}$
$+\int_{0}^{t}\Bigg{\{}e^{-H(\mu+\frac{1}{2})s-\frac{1}{2H}e^{-Hs}}+\left(2-e^{-Hs}\right)^{\mu+\frac{1}{2}}e^{-\frac{1}{2H}(2-e^{-Hs})}\Bigg{\}}\,ds\Bigg{]}\,.$
This indicates the rate of decay of the $L^{\infty}$-norm of solution for a large $t$.

### v2

@@
 Keywords: Klein-Gordon equation, Spherical waves, Expanding universe, Pionic atom
+MSC2020 Classification:: 35C15, 35Q40, 35Q75, 81T20
 School of Mathematical and Statistical Sciences,
@@
 1 Introduction. Main Results
-In this study, we produce an explicit wave function of the particles (fields) emitted to the FLRW universe with the scale factor generated by the de Sitter universe.
+In this study, we produce an explicit wave function of the particles (fields) emitted to the Friedmann-Lemaître-Robertson-Walker (FLRW) universe with the scale factor generated by the de Sitter space.
 We consider models of scalar fields with spherical symmetry at the initial time.
@@
 In the expanding universe, the particles obey the equation for the free particles.
-A purpose of this article is a derivation of the explicit form of exact solutions of the spherical waves in the curved spacetime.
-The models illustrate an instantaneous transition from the initial spacetime to the non-static FLRW universe with a de Sitter scale factor.
+The purpose of this article is to derive the explicit form of exact solutions of the spherical waves in curved spacetime.
+The models illustrate a transition from the initial spacetime to the non-static FLRW universe with a de Sitter scale factor.
 Thus, for the scalar field, we are interested in the Cauchy problem
-$\cases{\displaystyle\Box_{g}\Phi+\frac{m_{n}^{2}c^{2}}{\hbar^{2}}\Phi=0\,,\cr\displaystyle\Phi(x,0)=F(r,0)Y_{\ell m}(\theta,\phi)\,,\quad\Phi_{t}(x,0)=F_{t}(r,0)Y_{\ell m}(\theta,\phi)\,,}$
+$\cases{\displaystyle\Box_{g}\Phi-\frac{m_{n}^{2}c^{2}}{\hbar^{2}}\Phi=0\,,\cr\displaystyle\Phi(x,0)=F(r,0)Y_{\ell m}(\theta,\phi)\,,\quad\Phi_{t}(x,0)=F_{t}(r,0)Y_{\ell m}(\theta,\phi)\,,}$
 where $m_{n}$ is a mass, $F(r,t)$ is a time-dependent radial part of the initial wave, $F_{t}(r,t):={\partial_{t}F}(r,t)$, while $Y_{\ell m}(\theta,\phi)$ are spherical harmonics.
@@
 Consequently, the covariant d’Alembert’s operator in the spacetime with the metric of (1.1) is
-$\square_{g}\Phi=\frac{1}{\sqrt{|g|}}\frac{\partial}{\partial x^{i}}\left(\sqrt{|g|}g^{ik}\frac{\partial\Phi}{\partial x^{k}}\right).$
-We are interested in the explicit form of the spherical field in the de Sitter FLRW spacetime, which is a solution of the problem
+$\square_{g}\Phi=\frac{1}{\sqrt{|g|}}\frac{\partial}{\partial x^{\alpha}}\left(\sqrt{|g|}g^{\alpha\beta}\frac{\partial\Phi}{\partial x^{\beta}}\right),$
+where $\alpha$ and $\beta$ are running from 0 to $3$. We are interested in the explicit form of the spherical field in the de Sitter FLRW spacetime, which is a solution of the problem
 $\cases{\displaystyle\Phi_{tt}+3H\Phi_{t}-c^{2}e^{-2Ht}\Delta\Phi+\frac{m_{n}^{2}c^{4}}{\hbar^{2}}\Phi=0\,,\cr\displaystyle\Phi(x,0)=F_{0}(r)Y_{\ell m}(\theta,\phi)\,,\quad\Phi_{t}(x,0)=F_{1}(r)Y_{\ell m}(\theta,\phi)\,.}$ (1.2)
 Here, $Y_{lm}(\theta,{\phi})$ is a spherical harmonic, $\ell=0,1,\ldots$, $m=0,\pm 1,\pm 2\ldots$, $|m|\leq\ell$.
-In particular, the initial state can be given by a wave function $\mbox{\rm Exp}\left(-i\frac{E_{n}t}{\hbar}\right)\frac{R_{n\ell}(r)}{r}Y_{\ell m}(\theta,\phi)$ of the pionic atom (see, e.g., [8]).
+In particular, the initial data $\Phi(x,0)$ and $\Phi_{t}(x,0)$ can be given by a wave function of the pionic atom in Minkowski space, that is, by the functions $\frac{R_{n\ell}(r)}{r}Y_{\ell m}(\theta,\phi)$ and $\left(-i\frac{E_{n}}{\hbar}\right)\frac{R_{n\ell}(r)}{r}Y_{\ell m}(\theta,\phi)$, respectively (see, e.g., [2]).
 We will use a system of units in which the speed of light $c=1$.
@@
 We often omit the parameter $M$ in $K_{0}(z,t;M)$ and $K_{1}(z,t;M)$ when it does not lead to confusion.
-Denote $\phi(t):=(1-e^{-Ht})/H$.
+Denote $\phi(t):=(1-e^{-Ht})/H$ a conformal time.
+Note: the scalar curvature $R$ is encoded in these kernels via the Hubble constant $H=\sqrt{R/12}$.
 The main results of this study, which are obtained by the Integral Transform Approach (ITA) [6, 7], are the following two theorems.
@@
 $-\frac{1}{4}\ell(\ell+1)\frac{\tau}{r^{2}}\int_{r-\tau}^{r+\tau}(F_{0}(s)+F_{1}(s))F\left(1-\ell,\ell+2;2;\frac{\tau^{2}-(r-s)^{2}}{4rs}\right)ds\Bigg\}d\tau\,.$
+In the case of $m_{n}^{2}=2H^{2}\hbar^{2}$, the Klein-Gordon equation becomes the conformally invariant wave equation.
 It should be noted here that for negative integers $a$, the hypergeometric function $F(a,b;c;z)$ is a polynomial of degree a-a (see, e.g., [5]).
-In the next theorem, by imposing conditions on the behavior of the functions $F_{0}=F_{0}(r)$ and $F_{1}=F_{1}(r)$ for large $r$ one can appeal to the Hankel transform and write the representation of the field in the form that is more amendable for the $L^{2}({\mathbb{R}}^{3})$-estimates.
-In particular, this applies to the wave function of the pionic atom that instantly arrives in the de Sitter FLRW spacetime, because that wave function has an exponential decay at infinity of spatial coordinates (see, e.g., [8]).
+In the next theorem, by imposing conditions on the behavior of the functions $F_{0}=F_{0}(r)$ and $F_{1}=F_{1}(r)$ for large $r$ one can appeal to the Hankel transform and write the representation of the field in the form that is more amenable for the $L^{2}({\mathbb{R}}^{3})$-estimates.
+In particular, this applies to the wave function of the pionic atom that arrives in the de Sitter FLRW spacetime, because that wave function has an exponential decay at infinity of spatial coordinates (see, e.g., [2]).
 Theorem 1.2
@@
 $+\int_{0}^{\infty}\left(\int_{0}^{\infty}(F_{0}(\rho)+F_{1}(\rho))J_{\ell+\frac{1}{2}}(\lambda\rho)\,\rho^{3/2}d\rho\right)\sin\left(\lambda\phi(t)\right)J_{\ell+\frac{1}{2}}(r\lambda)\,d\lambda\Bigg\}\,.$
+In Corollary 3.2, we prove the exponential decay of the solution’s tail in the case where the radial part of the initial function is generated by the wave function of the pionic atom.
+Moreover, it is also shown that the obtained decay is optimal.
 Below, we provide some examples of such an initial state $(\Phi(x,0),\Phi_{t}(x,0))$ of the field.
@@
 A pion is an example of a particle with electromagnetic interactions that obeys the Klein-Gordon equation.
-In fact, by pionic atom, one can test the validity of the Klein-Gordon equation (see, e.g., [10, Sec.18.4]).
-In (1.2), the pion has instantly left the atom, the Coulomb field is no longer in action, and spacetime has instantly become inflationary.
+In fact, by pionic atom, one can test the validity of the Klein-Gordon equation (see, e.g., [8, Sec.18.4]).
+In (1.2), the pion has left the atom, the Coulomb field is no longer in action, and spacetime has become inflationary.
 Thus, the free pion propagates in the expanding universe and carries information about its previous state.
-Another possible application can be to the problem of the existence of quasinormal modes on de Sitter space (see [12]).
-Although spherically symmetric solutions for the Klein-Gordon and Dirac equations in the Minkowski spacetime have been known for a long time (see, e.g., [4, §24], [7], [15], and references therein), explicit exact formulas in the curved spacetimes are known only for a few special cases (see, e.g., [2, 6, 18, 25]).
+Although spherically symmetric solutions for the Klein-Gordon and Dirac equations in the Minkowski spacetime have been known for a long time (see, e.g., [11, §24], [12], [13], and references therein), explicit exact formulas in the curved spacetimes are known only for a few special cases (see, e.g., [14, 15, 16, 17]).
 We believe that the obtained explicit formulas of Theorem 1.2 can be useful, in particular, in the study of caustics for spherical waves in curved spacetime (see, e.g., [12]).
 In the forthcoming paper, the representation formulas of Theorems 1.1, 1.2 will be used for the derivation of the exact spherical solutions to the spin-$\frac{1}{2}$ fields in the expanding universe.
+Another possible application of our formulas can be to the problem of the existence of quasinormal modes (QNMs) on a non-static FLRW universe with a de Sitter scale factor.
+The existence of QNMs on de Sitter spacetime has been much studied in the literature; a detailed analysis of the literature on the spherical waves and QNMs has been given in [18].
+In [18], it has been demonstrated that QNMs exist for massive scalar fields on static de Sitter space in all spacetime dimensions greater than one and for all complex-valued scalar field masses.
+The Taylor expansion with the recursion relation for the coefficients has been presented for calculating of QNMs and their corresponding mode solutions.
+The proof of the convergence of the series has been done for the analytic data.
+The study of the spherically symmetric solutions of the wave equation in the de Sitter space in the Bondi coordinates was carried out in [1].
+In those coordinates, the problem was reduced to an integro-differential equation with the initial data on a light cone.
+The main results state existence and uniqueness, as well as decay of the solution in the Bondi coordinates and its limit at infinity.
 The rest of this paper is organized as follows.
@@
 We look for the solution $F=F(r,t)$ of equation (2.3) satisfying (2.4).
-For example, if the initial state $F_{0}(r)$ is the the wave function of the pionic atom [8, pp.188-189], [10, Sec.18.4], and [14, §26, Sec 2] in the Minkowski space, then
+For example, if $F_{0}(r)$ is generated by the wave function of the pionic atom [2, pp.188-189], [8, Sec.18.4], and [3, §26, Sec 2] in the Minkowski space, then
 $F_{0}(r)=C_{n,\ell,\mu}r^{\mu-\frac{1}{2}}e^{-r/2}L_{n-\ell-1}^{2\mu}(r),\quad L_{n-\ell-1}^{2\mu}(r)\quad\mbox{\rm is Laguerre polynomial},\,\,r>0,$ (2.5)
@@
 $\mu=\sqrt{\left(\ell+\frac{1}{2}\right)^{2}-Z^{2}\alpha^{2}}\geq 0\,.$
-If $\ell=0$, while $Z=1$ and $Z=2$, then $\mu-\frac{1}{2}=-0.0000532822$ and $\mu-\frac{1}{2}=-0.000213163$, respectively.
+If $\ell=0$, while the atomic number $Z=1$ and $Z=2$, then $\mu-\frac{1}{2}=-0.0000532822$ and $\mu-\frac{1}{2}=-0.000213163$, respectively.
 For the case of $F_{1}(r)=0$, assume that $F_{0}(r)$ is smooth on ${\mathbb{R}}\setminus\{0\}$. Define the operator
@@
 $\widetilde{F}_{0}^{(\ell)}(r)=r^{\ell}\int_{0}^{r}s^{-\ell}F_{0}^{(\ell+1)}(s)\,ds\,.$ (2.8)
+Here $a_{k,\alpha,\beta}^{(\ell)}$ and $b_{k,\alpha,\beta}^{(\ell)}$ are some suitable coefficients.
 For the first three values of $\ell$, if $F_{0}$ is given by (2.5), then the function $F(r,t)$ can be a distributional solution from the Sobolev space.
@@
 $\leq$ $C_{n,\ell,\mu}\int_{t-r}^{t+r}e^{-\tau/2}\tau^{a}d\tau\,$,
-where $a:=\mu-\frac{1}{2}+n-\ell-1+k$. The last integral can be evaluated in terms of the incomplete gamma function ([3, Ch.9]):
+where $a:=\mu-\frac{1}{2}+n-\ell-1+k$. The last integral can be evaluated in terms of the incomplete gamma function ([5, Ch.9]):
 $\left|\int_{r-t}^{r+t}F_{0}(\tau)\tau^{k}d\tau\right|$ $\leq$ $C_{n,\ell,\mu}2^{a+1}\left|\Gamma\left(a+1,\frac{t-r}{2}\right)-\Gamma\left(a+1,\frac{r+t}{2}\right)\right|$
@@
 $F_{\nu}(s):=\int_{0}^{\infty}rf(r)J_{\nu}(sr)\,dr\,.$
-with $\nu>-1/2$ ([16, Sec.11.5.1], [19, Sec.14.3], [27, Sec.14.3]). Let $f(r)$ be a function defined for $r>0$. The inversion formula yields
+with $\nu>-1/2$ ([24, Sec.11.5.1], [25, Sec.14.3], [26, Sec.14.3]). Let $f(r)$ be a function defined for $r>0$. The inversion formula yields
 $f(r)=\int_{0}^{\infty}sF_{\nu}(s)J_{\nu}(sr)\,ds\,.$
@@
 The spherical Klein-Gordon field in the Minkowski space is given by
-$\Phi(r,\theta,\phi,t)$ $=$ $Y_{\ell m}(\theta,\phi)\frac{1}{\sqrt{r}}\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{0}(\rho)J_{\ell+\frac{1}{2}}(\rho\lambda)\rho^{3/2}\,d\rho\right)\cos\left(\lambda t\right)J_{\ell+\frac{1}{2}}(r\lambda)\lambda\,d\lambda$
-$-Y_{\ell m}(\theta,\phi)\frac{1}{\sqrt{r}}\int_{0}^{t}\Bigg{\{}\frac{m_{0}t}{\sqrt{t^{2}-\tau^{2}}}J_{1}\left(m_{0}\sqrt{t^{2}-\tau^{2}}\right)-J_{0}\left(m_{0}\sqrt{t^{2}-\tau^{2}}\right)\Bigg{\}}$
-$\times\Bigg{\{}\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{1}(\rho)J_{\ell+\frac{1}{2}}(\rho\lambda)\rho^{3/2}\,d\rho\right)\cos\left(\lambda\tau\right)J_{\ell+\frac{1}{2}}(r\lambda)\lambda\,d\lambda\Bigg{\}}\,d\tau$
+$\Phi(r,\theta,\phi,t)$ $=$ $Y_{\ell m}(\theta,\phi)\frac{1}{r}\Bigg\{\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{0}(\rho)J_{\ell+\frac{1}{2}}(\lambda\rho)\,\rho^{2}d\rho\right)\cos(\lambda t)J_{\ell+\frac{1}{2}}(r\lambda)\,\lambda d\lambda$
+$-\int_{0}^{t}\frac{mt}{\sqrt{t^{2}-\tau^{2}}}J_{1}\left(m\sqrt{t^{2}-\tau^{2}}\right)\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{0}(\rho)J_{\ell+\frac{1}{2}}(\lambda\rho)\,\rho^{2}d\rho\right)\cos(\lambda\tau)J_{\ell+\frac{1}{2}}(r\lambda)\,\lambda d\lambda\,d\tau$
+$+\int_{0}^{t}J_{0}\left(m\sqrt{t^{2}-\tau^{2}}\right)\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{1}(\rho)J_{\ell+\frac{1}{2}}(\lambda\rho)\,\rho^{2}d\rho\right)\cos(\lambda\tau)J_{\ell+\frac{1}{2}}(r\lambda)\,\lambda d\lambda\,d\tau\Bigg\}$
 for all $x\in{\mathbb{R}}^{3}$, $t>0$.
@@
 (For the definition of the Huygens’ principle see, e.g., [28].)
+Thus, one can regard ITA as an analytical mechanism that, from the massless virtual field in the Minkowski space, generates a massive field in the curved spacetime.
+The integral transform condenses the evolution of that virtual scalar wave in the conformal time from its entry into the expanding universe to the moment of observation.
 3.1 Proof of Theorem 1.1
@@
 $\left|\Phi(r,\theta,\phi,t)-e^{-Ht}v_{F_{0}Y_{\ell m}}(r,\theta,\phi,\phi(t))\right|$
-$\leq$ $C(r)e^{-\frac{3}{2}Ht}\left|Y_{\ell m}(\theta,\phi)\right|\cases{e^{\frac{1}{2}Ht}\quad\mbox{\rm if}\quad 3H\hbar/2>m_{n}>\sqrt{2}H\hbar\,,\cr e^{t\sqrt{\frac{9}{4}H^{2}-\frac{m_{n}^{2}}{\hbar^{2}}}}\quad\mbox{\rm if}\quad m_{n}<\sqrt{2}H\hbar\,.}$
-If $m_{n}=\sqrt{2}H\hbar$, then
-$\left|\psi(x,t)-e^{-Ht}v_{\psi_{0}}(x,\phi(t))\right|$ $\leq$ $C(r)e^{-Ht}\left|Y_{\ell m}(\theta,\phi)\right|\,.$
+$\leq$ $C(r)e^{-\frac{3}{2}Ht}\left|Y_{\ell m}(\theta,\phi)\right|\cases{e^{\frac{1}{2}Ht}\quad\mbox{\rm if}\quad 3H\hbar/2>m_{n}\geq\sqrt{2}H\hbar\,,\cr e^{t\sqrt{\frac{9}{4}H^{2}-\frac{m_{n}^{2}}{\hbar^{2}}}}\quad\mbox{\rm if}\quad m_{n}<\sqrt{2}H\hbar\,.}$
 Here $v_{F_{0}Y_{\ell m}}(x,\phi(t))$ is given by Lemma 2.4,
@@
 This indicates the rate of decay of the $L^{\infty}$-norm of solution for a large $t$.
+Acknowledgments
+The author is indebted to the reviewers for the valuable remarks and suggestions, which improved the text.